OpenStudy (anonymous):

simplify (2x^3y^-1)^2 (-4xy^3)2

OpenStudy (zehanz):

If you apply the rules \[(ab)^c=a^cb^c\]\[(a^b)^c=a^{bc}\]\[a^ba^c=a^{b+c}\]you're done! So let me begin:\[(2x^3y^{-1})^2 \cdot (-4xy^3)^2=2^2(x^3)^2(y^{-1})^2 \cdot (-4)^2x^2(y^3)^2\]I applied the first rule. Now you apply the second!

OpenStudy (zehanz):

Have you got the right answer in the mean time?